Assume all future cash flows (FCFF, FCFE, and Dividends) will grow at a constant sustainable growth rate (g) in perpetuity; note this growth rate is dependent on a companys return on equity and dividend payout policy. The companys cost of debt is 20% while the companys cost of equity is 30%; the company has an effective tax rate of 35%. Use the following information in the table below to help answer problems 26-27: FCFE Today (T O) FCFF Today (1-0) Shareholder Equity in millions 800 750 Total Assets Net Income Dividends Shares Outstanding 030 1000 100 40

27. A private equity group who wants to purchase all of the company’s assets would be willing to pay approximately __________. Do not use the DDM. Use either FCFF or FCFE, whichever is appropriate.

  1. $9,583M

  2. $10,222M

  3. $10,518M

  4. $17,969M

  5. $19,167M
Answer #1

Cost of capital WACC=600/1000*20%*(1-35%)+400/1000*30%=19.8%

Enterprise Value=750*(1+15%)/(19.8%-15%)=17969 M

Option D